# Generalized Bayesian deep reinforcement learning

### v1

@@ 2.1 Expected Thompson sampling
While updating the posterior at each interaction round can be computationally expensive, we mitigate this by updating the posterior after a set amount of agent-environment interaction. This makes the learning process more efficient without sacrificing performance.
In reinforcement learning (RL) tasks, terminal states often signify the end of an episode, which restarts from an initial state. These episodes, each beginning at an initial state and ending at a terminal state, are mutually independent. Tasks with clear terminal conditions, such as games that end with a win or loss, are referred to as episodic tasks. In such tasks, it is standard to update the policy at the end of each episode and use the updated policy in future interactions.
In contrast, infinite-horizon tasks have no defined terminal state. Here, a custom interaction length can be chosen to determine when to update the policy. Whether the task is episodic or not, we treat each interaction sequence as an episode and, for simplicity, assume a constant episode length $\tau$ in our analysis. However, our theorems can easily be extended to handle varying episode lengths. Next, we describe our episodic policy update strategy using ETS.
In settings where the action space $\mathcal{A}$ is finite and discrete, the optimal policy is found by optimizing the $Q$-function. In model-free approaches, data from real interactions with the environment is used for $Q$-function estimation. However, in Thompson-sampling-based methods (Dimitrakakis and Tziortziotis, 2013), a sample is drawn from the posterior of the model parameters to simulate longer trajectories (which means more data) of interactions. The $Q$-function is then estimated based on these simulated trajectories.
In the ETS framework, we extend this idea by drawing multiple samples from the posterior distribution. These posterior samples are used to simulate multiple interaction trajectories, leading to a more accurate estimate of the $Q$-function. If the posterior distribution of the model parameters after the $k$th episode is $\pi_{k}(\theta)=\pi(\theta|H_{\tau k})$, and $\bm{\theta}^{(k)}=\{\theta^{ki}\}_{i=1}^{n}$ represents a set of $n$ posterior samples, then the $Q$-function for policy $\mu$, based on the simulated trajectories, is estimated as:
@@
ETS can be easily integrated with policy iteration methods by performing the policy evaluation using equation (3). In the following, we provide pseudocode for a policy iteration method incorporating ETS.
@@
Generate $H_{\tau}=\{(s_{t},a_{t},r_{t+1},s_{t+1})\}_{t=1}^{\tau}$ via random policy ($=\mu^{(1)}$, say )
@@
     Sample $\bm{\theta}^{(k)}=\{\theta^{ki}\}_{i=1}^{n}\sim\pi_{k}(\cdot)$
@@
The chain task is a very simple task which has the typical exploration-exploitation dilemma. Although there is a small reward at the first state, there is a bigger reward for reaching the last state. For a fairly big horizon, taking the right action is the optimal policy which we want our algorithm to learn.
Note that, here, the true model for the environment’s dynamics can be expressed with $Binomial(p_{0}=0.8)$ transition probability and rewards distributed as $\mathcal{N}(\overline{r}_{0},I_{5})$, with $\overline{r}_{0}=(0.2,0,0,0,1)$ for the 5 states. Here we assume the parametric model ($\mathbb{M}_{\theta}$) for the underlying MDP is known to us except for the true model parameters, $\theta_{0}=(p_{0},\overline{r}_{0})$. Since the true underlying MDP $\mathbb{M}_{\theta_{0}}$ belongs to the class $\mathbb{M}_{\theta}$, this is an example of well-specified model. To infer the model parameters, we started with assigning a Beta prior on the transition probabilities and a Gaussian prior on the mean rewards. Using the conjugacy of the prior distributions, we drew samples from the exact conjugate posterior distribution to perform ETS.
For policy learning, we have used a dynamic programming algorithm called Backwards Induction (BI) following the implementation of Dimitrakakis and Ortner (2022). This method is well-suited for a finite-horizon RL task with finite state and action space. The BI algorithm uses Bellman’s equations to derive the action value function starting from the terminal state and going backwards to calculate it for each state recursively. To accommodate ETS, within each episode of interaction, for each state, we calculate the $Q$ function based on the all the sampled MDPs according to equation (3) and choose the greedy action that maximizes the pooled estimate of the $Q$ function.
Now let $\mu_{0}$ be the optimal policy under the true MDP which is unknown to the agent. Suppose the agent interacts with the environment over epochs or episodes, $k=1,2,\ldots$ and $\hat{\mu}_{k}$ is the policy to determine action from a given state during the $k$th episode using ETS. The total regret of ETS due to using the sequence of policies $\hat{\mu_{1}},\hat{\mu_{2}},\dots$ over a total of $T$ timesteps is defined as the sum of episodic regrets as shown below,
@@
(a) Plot of total regret for different posterior sample size
(b) Plot of episodic regrets over the episodes
@@
Bayesian model-based RL requires a Bayesian treatment of the model parameters. However, when dealing with complex models with intractable likelihoods (e.g. when the analytical form for $P_{\theta_{st}}$ and $P_{\theta_{r}}$ is unavailable for a parametric MDP), directly calculating the posterior distribution becomes challenging. Nevertheless, for most of the assumed parametric MDP models we can directly sample from them, a distinctive feature that we capitalize upon. This capability to derive samples directly from the model suggests using a likelihood-free inference framework. This approach entails generating data by running simulations based on the model parameters. The samples we generate from the simulator contain a spectrum of potential outcomes under the model, given the population parameters. By using methods available under the broad umbrella of likelihood-free inference (LFI), these samples let us make inferences without having to confront the mathematical intractability of the likelihood function. Two of the most popular classes of LFI methods are approximate Bayesian computation (ABC) (Lintusaari et al., 2017) and Bayesian synthetic likelihood (BSL) (Price et al., 2018), of which ABC has been used before for Thompson Sampling in (Dimitrakakis and Tziortziotis, 2013). They approximate the intractable likelihood function either implicitly or explicitly using these samples. The asymptotic contraction of these approximate posteriors towards the true parameter value depends upon the choice of the summary statistics and some conditions being satisfied by the chosen summary statistics (Frazier et al., 2018; Li and Fearnhead, 2018; Frazier et al., 2023), which are difficult to verify in practice.
@@ 3.1 Scoring rule posterior
In the setup of LFI, as the likelihood function $p(\mathbf{y}^{obs}\mid\theta)$ for a model with observed data $\mathbf{y}^{obs}$ is intractable, we are unable to obtain the posterior distribution using Bayes Theorem
@@
when parameters $\theta$ and $\pi(\theta)$ are the parameters and prior distribution on our parameters. To mitigate this, Pacchiardi et al. (2024b) consider loss functions $S(P,\mathbf{y})$ known as scoring rules (Gneiting and Raftery, 2007) that measure the fit between the distribution $P_{\theta}$ of the data under parameter $\theta$, and an observed data point $\mathbf{y}$.
The scoring rule posterior is then defined as follows:
$\pi_{S}\left(\theta\mid\mathbf{y}^{obs}\right)\propto\pi(\theta)\exp\left\{-wS\left(P_{\theta},\mathbf{y}^{obs}\right)\right\}$
Comparing the two expressions, we note that the (negative) log-likelihood function can itself be considered a scoring rule (known as log-score, Dawid and Musio, 2014) and that we have introduced an additional parameter $w$, which is known as the learning rate in generalized Bayesian inference (Holmes and Walker, 2017) controlling the relative weighting of the observations relative to the prior. For this study we keep $w$ fixed at $1$. However, $w$ could be further tuned to adapt the contraction rate of the scoring rule posterior (Bissiri et al., 2016).
A scoring rule is said to be proper if its expected value with respect to the observed data $\mathbf{y}^{obs}$ is minimized when the assumed distribution $P_{\theta}$ is equal to the distribution $P_{0}$ generating the observed data $\mathbf{y}^{obs}$. If $P_{\theta}=P_{0}$ is the unique minimum, the scoring rule is said to be strictly proper. A strictly proper scoring rule is desired as this property further ensures that the best prediction one can make, according to the scoring rule, is exactly the truth. The existence of a unique minimum, i.e., the true distribution, provides a strong motivation for models to accurately capture the underlying distribution of the data. Theoretical properties of asymptotic normality and generalization bound for scoring rule posteriors when the scoring rule is strictly proper have been studied in Giummolè et al. (2019); Pacchiardi et al. (2024b). Additionally, the scoring rule posterior for some scoring rules (e.g. energy score or kernel score) exhibits robustness against outliers compared to the standard Bayes posterior using log-score.
@@
$S_{E}^{(\beta)}(P_{\theta},\mathbf{y})=2\cdot\mathbb{E}||X-\mathbf{y}||_{2}^{\beta}-\mathbb{E}||X-X^{\prime}||_{2}^{\beta},\quad X,X^{\prime}\sim P_{\theta};\quad\beta\in(0,2)$ (4)
This is a strictly proper scoring rule for the class of probability measures $\mathcal{P_{\theta}}$ for which $\mathbb{E}_{X\sim P_{\theta}}||X||^{\beta}<\infty$ for all $P_{\theta}\in\mathcal{P}$ (Gneiting and Raftery, 2007). The related divergence is the square of the energy distance, which is a metric between probability distributions. Further the energy score can be unbiasedly estimated using $\mathbf{x}_{j}\sim P,j=1,\ldots,m$ which are independent and identically distributed samples from the model $P$. The unbiased estimate for the energy score can be obtained by Monte Carlo estimates of the expectations in $S_{\mathrm{E}}^{(\beta)}(P,y)$
$\hat{S}_{\mathrm{E}}^{(\beta)}\left(\left\{x_{j}\right\}_{j=1}^{m},\mathbf{y}\right)=\frac{2}{m}\sum_{j=1}^{m}\left\|x_{j}-\mathbf{y}\right\|_{2}^{\beta}-\frac{1}{m(m-1)}\sum_{\begin{subarray}{c}j,k=1\\ k\neq j\end{subarray}}^{m}\left\|x_{j}-x_{k}\right\|_{2}^{\beta}\quad,\beta\in(0,2).$ (5)
For our implementation, we will consider $\beta=1$ and will write $S_{\mathrm{E}}^{(1)}(P,\mathbf{y})$ simply as $S_{\mathrm{E}}(P,\mathbf{y})$.
Note that, SRs such as energy scores can be expressed as $S(P_{\theta},\mathbf{y})=\mathbb{E}_{X,X^{\prime}\sim P_{\theta}}g(X,X^{\prime},\mathbf{y})$. Moreover, suppose the simulator model $P_{\theta}$ can also be represented as $X=h_{\theta}(Z)$, with $Z\sim Q$ where, the distribution $Q$ is independent of $\theta$. It has been shown in Pacchiardi et al. (2024b) that, if both $g$ and $h_{\theta}$ differentiable then an interchange of expectation and gradient step enables unbiased estimation of the gradient of the SR $S(P_{\theta},\mathbf{y})$ by using some random draws of $z_{i}\sim Q$ for $i=1,2,\ldots m$. This property is particularly useful in high-dimensional parameter spaces, where efficient exploration requires gradient information.
@@ 3.2 Prequential scoring rule posterior
As mentioned before, for a complex MDP we may not have the analytical form for the conditional probabilities $P_{\theta_{st}}$ and $P_{\theta_{r}}$ used in equation (1) to define the loglikelihood function. Hence combining the idea of scoring posterior explained in the previous section with the idea of prequential (predictive sequential) SR proposed in Pacchiardi et al. (2024a), here we propose prequential scoring rule posterior. For prequential SR, the original SRs are used to assess the goodness of fit of the one step conditional distributions using the observations at the next step given the observations at the previous step. Hence the prequential SR for a parametric MDP ($\mathbb{M}_{\theta}$) based on $H_{T}$ using a scoring rule $S$ would be given by,
@@
When the scoring rule $S$ is (strictly) proper then the prequential scoring rule $\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})$ is also (strictly) proper for the Markovian conditional distributions over the next state and reward given the previous state-action pair as shown in Theorem 2 of Pacchiardi et al. (2024a). If the true MDP belongs in class of MDPs =$\mathcal{M}=\{\mathbb{M}_{\theta}:\theta\in\Theta\}$ with the true parameter value being $\theta_{0}$, then the minimizer of the expected prequential scoring rule,
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{H_{T}\sim\mathbb{M}_{0}}\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})$ (7)
would coincide with the true parameter $\theta_{0}$.
We notice, if we have access to the analytical form of $P_{\theta_{st}}$ and $P_{\theta_{r}}$, then using negative log-score the above equation reduces to equation (1).
If we can estimate the SR $S$ as shown in equation (5) unbiasedly, then the prequential SR $\operatorname{\mathcal{PS}}$ can also be estimated unbiasedly by generating $m$ simulations of next states and reward until time $T$, conditioned on previously observed state and action. Further in many MDPs, the reward distribution is chosen deterministically based on the problem and it is considered as a constant function of state and action. Hence, without loss of generality, in those cases, our $\operatorname{\mathcal{PS}}$ simplifies into
$\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})=\sum_{t=1}^{T}S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1}),$ (8)
and $\theta$ only contains parameters $\theta_{st}$. Based on this, we can define our prequential scoring rule posterior on the parameter $\theta$ of the parametric MDP as,
$\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid H_{T}\right)\propto\pi(\theta)\exp\left\{-w\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})\right\},$ (9)
when $\pi(\theta)$ is the prior distribution on our parameters. Note that choosing values of $w$ other than $1$ can be interpreted as a form of annealing (heating it for $w<1$ and cooling it for $w>1$, we have taken $w=1$ for our experiments) applied to the target posterior.
Asymptotic properties of the prequential scoring rule posterior
Under the assumption of finite action space ($\mathcal{A}$ is countable and $|\mathcal{A}|=n_{a}<\infty$), we can also define prequential scoring rules for each of the actions as following,
$\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})=\sum_{t=1}^{T}S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1})\mathbb{I}_{a_{t}=a},$
where $\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})=\sum_{a\in\mathcal{A}}\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})$. Based on this, we can rewrite our prequential scoring rule posterior on the parameter $\theta$ of the parametric MDP as,
$\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid H_{T}\right)\propto\pi(\theta)\exp\left\{-w\sum_{a\in\mathcal{A}}\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})\right\},$ (10)
when $\pi(\theta)$ is the prior distribution on our parameters. Further we denote, $T_{a}$ as the set consisting of timepoints when action $a$ was taken until time $T$, i.e., $T_{a}=\{t\leq T:a_{t}=a\}$ for each $a\in\mathcal{A}$. Assuming mild conditions on each of the action-specific Prequential SR loss functions $\operatorname{\mathcal{PS}}^{a}$ for $a\in\mathcal{A}$, the Prequential SR posterior satisfies a Bernstein-von Mises (BvM) theorem for finite action spaces. This theorem ensures the asymptotic normality of the corresponding posterior, with results analogous to the asymptotic behavior of the standard Bayesian posterior.
@@
Next, we will prove the BvM theorem under some assumptions. Before stating the assumptions, we first clarify some mathematical notations. We assume $\Theta\subseteq\mathbb{R}^{d}$ and the minimizer of the prequential scoring rule $\theta^{*}\in E$, where $E\subseteq\Theta$ is open (in $\mathbb{R}^{d}$ ) and bounded. The prior $\pi:\Theta\rightarrow\mathbb{R}$ is a probability density function with respect to Lebesgue measure such that $\pi$ is continuous at $\theta^{*}$ and $\pi\left(\theta^{*}\right)>0$. For the sake of notational simplicity, from now on we will use $\operatorname{\mathcal{PS}}_{T}(\theta)$ to denote the prequential scoring rule $\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})$ between the MDP $\mathbb{M}_{\theta}$ and $H_{T}$ and similarly $\operatorname{\mathcal{PS}}^{a}_{T}(\theta)$ to denote $\operatorname{\mathcal{PS}}^{a}(\mathbb{M}_{\theta},H_{T})$ for each action $a$. We denote the gradient, the matrix of second order derivatives and the tensor of third order derivatives of a function $f(\theta)$ with respect to $\theta$ as $f^{\prime}(\theta)=\left(\frac{\partial f(\theta)}{\partial\theta^{i}}(\theta)\right)_{i=1}^{d}\in\mathbb{R}^{d}$, $f^{\prime\prime}(\theta)=\left(\frac{\partial f}{\partial\theta^{i}\partial\theta^{j}}(\theta)\right)_{i,j=1}^{d}\in\mathbb{R}^{d^{2}}$ and $f^{\prime\prime\prime}(\theta)=\left(\frac{\partial f}{\partial\theta^{i}\partial\theta^{j}\partial\theta_{k}}(\theta)\right)_{i,j,k=1}^{d}\in\mathbb{R}^{d^{3}}$ respectively. For a given $\theta\in\mathbb{R}^{d}$ and $r>0$, we would denote the open ball of radius $r$ around $\theta$ as $B_{r}(\theta)=\{\theta^{\prime}\in\mathbb{R}^{d}:|\theta^{\prime}-\theta|<r\}$, where ||$\cdot$|| ⋅ | stands for Euclidean norm.
We first prove Lemma 1 under the following assumptions, which guarantee the existence of a sequence of consistent estimators of the true model parameter that converges to the expected SR minimizer defined in equation (7). It is important to note that Assumption A3 is a rather strong assumption regarding the convexity of each of the action specific prequential SRs. However, it is sufficient to verify Assumption A2 in conjunction with Assumption A1 to show the consistency.

### v2

@@ 2.1 Expected Thompson sampling
To reduce computational overhead, we update the posterior after a fixed number of interactions rather than after every step. In RL, episodic tasks—with clear terminal states—naturally allow for policy updates at the episode end. For non-episodic (infinite-horizon) tasks, we define episodes of fixed length $\tau$ and update the policy accordingly. While we assume constant episode length for simplicity, our results can be extended to variable lengths. We now outline our episodic policy update strategy using ETS.
In settings with a finite, discrete action space $\mathcal{A}$, the optimal policy is obtained by maximizing the $Q$-function. Let, $\pi_{k}(\theta)=\pi(\theta\mid h_{\tau k})$ denote the posterior distribution over model parameters after the $k$-th episode. In TS-based methods (Dimitrakakis and Tziortziotis, 2013), a single parameter sample $\theta_{k}\sim\pi_{k}$ is drawn, and the policy $\mu$ is optimized based on the estimated $Q$-function $Q_{\mu}^{\theta_{k}}$ from simulated trajectories using the model $\mathbb{M}_{\theta_{k}}$.
In contrast, we propose to use multiple posterior samples. Let $\bm{\theta}^{(k)}=\{\theta_{ki}\}_{i=1}^{n}$ represents a set of $n$ posterior samples from $\pi_{k}$. The $Q$-function for policy $\mu$ under ETS is then estimated by:
@@
ETS can be easily integrated with policy iteration methods by performing the policy evaluation using equation (3). In the following, we provide pseudocode as Algorithm 1.
@@
Observe $h_{\tau}=\{(s_{t},a_{t},r_{t+1},s_{t+1})\}_{t=1}^{\tau}$ by playing a random policy ($=\mu^{(1)}$, say )
@@
     Sample $\bm{\theta}^{(k)}=\{\theta_{ki}\}_{i=1}^{n}\sim\pi_{k}(\cdot)$
@@
The chain task is a very simple task that has the typical exploration-exploitation dilemma. Although there is a small reward at the first state, there is a bigger reward for reaching the last state. For a fairly big horizon, taking the right action is the optimal policy that we want our algorithm to learn.
Note that, here, the true model for the environment’s dynamics can be expressed with $Binomial(p_{0}=0.8)$ transition probability and rewards distributed as $\mathcal{N}(\overline{r}_{0},I_{5})$, with $\overline{r}_{0}=(0.2,0,0,0,1)$ for the 5 states. Here we assume the parametric model ($\mathbb{M}_{\theta}$) for the underlying MDP is known to us except for the true model parameters, $\theta_{0}=(p_{0},\overline{r}_{0})$. Since the true underlying MDP $\mathbb{M}_{0}=\mathbb{M}_{\theta_{0}}$ belongs to the class $\mathbb{M}_{\theta}$, this is an example of well-specified model. To infer the model parameters, we started by assigning a Beta prior to the transition probabilities and a Gaussian prior to the mean rewards. Using the conjugacy of the prior distributions, we drew samples from the exact conjugate posterior distribution to perform ETS.
For policy learning, we have used a dynamic programming algorithm called backward induction (BI) following the implementation of Dimitrakakis and Ortner (2022). This method is well-suited for a finite-horizon RL task with finite state and action space. The BI algorithm uses Bellman’s equations to derive the action value function starting from the terminal state and going backward to calculate it for each state recursively. To accommodate ETS, within each episode of interaction, for each state, we calculate the $Q$ function based on all the sampled MDPs according to equation (3) and choose the greedy action that maximizes the pooled estimate of the $Q$ function.
(a) Plot of total regret for different posterior sample size
(b) Plot of episodic regrets over the episodes
Now let $\mu_{0}$ be the optimal policy under the true MDP which is unknown to the agent. Suppose the agent interacts with the environment over epochs or episodes, $k=1,2,\ldots$ and ${\mu}^{(k)}$ is the policy to determine action from a given state during the $k$th episode using ETS. The total regret of ETS due to using the sequence of policies ${\mu}^{(1)},{\mu}^{(2)},\dots$ over a total of $T$ timesteps is defined as the sum of episodic regrets as shown below,
@@
Bayesian model-based RL requires a Bayesian treatment of the model parameters. However, when dealing with complex models with intractable likelihoods (e.g. when the analytical form for $P_{\theta_{st}}$ and $P_{\theta_{r}}$ is unavailable for a parametric MDP), directly calculating the posterior distribution becomes challenging. Nevertheless, for most of the assumed parametric MDP models, we can directly sample from them, a distinctive feature that we capitalize upon. This capability to derive samples directly from the model suggests using a likelihood-free inference framework. This approach entails generating data by running simulations based on the model parameters. The samples we generate from the simulator contain a spectrum of potential outcomes under the model, given the population parameters. By using methods available under the broad umbrella of likelihood-free inference (LFI), these samples let us make inferences without having to confront the mathematical intractability of the likelihood function. Two of the most popular classes of LFI methods are approximate Bayesian computation (ABC) (Lintusaari et al., 2017) and Bayesian synthetic likelihood (BSL) (Price et al., 2018), of which ABC has been used before for Thompson Sampling in (Dimitrakakis and Tziortziotis, 2013). They approximate the intractable likelihood function either implicitly or explicitly using these samples. The asymptotic contraction of these approximate posteriors towards the true parameter value depends upon the choice of the summary statistics and some conditions being satisfied by the chosen summary statistics (Frazier et al., 2018; Li and Fearnhead, 2018; Frazier et al., 2023), which are difficult to verify in practice.
@@ 3.1 Scoring rule posterior
For models with intractable likelihood function $p(\mathbf{y}^{obs}\mid\theta)$, the posterior distribution cannot be computed directly via Bayes’ theorem:
@@
where $\theta$ is the model parameter and $\pi(\theta)$ is a prior distribution on the parameter space $\Theta$. To mitigate this, Pacchiardi et al. (2024b) consider loss functions $S(P_{\theta},\mathbf{y})$ known as scoring rules (Gneiting and Raftery, 2007) that measure the fit between the distribution $P_{\theta}$ of the data under parameter $\theta$, and an observed data point $\mathbf{y}$. These loss functions allow inference without access to the likelihood, requiring only the ability to simulate from $P_{\theta}$.
The likelihood-free scoring rule posterior is then defined as follows:
$\pi_{S}\left(\theta\mid\mathbf{y}^{obs}\right)\propto\pi(\theta)\exp\left\{-wS\left(P_{\theta},\mathbf{y}^{obs}\right)\right\}.$
Comparing the two expressions, we note that the (negative) log-likelihood function can itself be considered a scoring rule (known as log-score, (Dawid and Musio, 2014)) and that we have introduced an additional parameter $w$, which is known as the learning rate in generalized Bayesian inference (Holmes and Walker, 2017) controlling the relative weighting of the observations relative to the prior. Next, we discuss some properties of different types of scoring rules that ensure asymptotic contraction of the resulting posterior distributions.
Suppose $P_{0}$ is the underlying true data-generating process. Then, we can define the expected scoring rule as, $\tilde{S}(P_{\theta},P_{0})=\mathbb{E}_{Y\sim P_{0}}S(P_{\theta},Y)$. A scoring rule $S$ is said to be proper if $\tilde{S}(P_{\theta},P_{0})$ is minimized when the assumed distribution $P_{\theta}$ is equal to the distribution $P_{0}$ generating the observed data $\mathbf{y}^{obs}$. If $P_{\theta}=P_{0}$ is the unique minimum, the scoring rule is said to be strictly proper. Thus, when $S$ is strictly proper we can define a statistical divergence between the distributions $P_{\theta}$ and $P_{0}$ as:
$D(P_{\theta},P_{0})=\tilde{S}(P_{\theta},P_{0})-\tilde{S}(P_{0},P_{0})$ (4)
This divergence is non-negative and equals zero if and only if $P_{\theta}=P_{0}$. This is a generalized divergence that measures how well the model $P_{\theta}$ approximates the true data-generating process $P_{0}$. If we take the scoring rule to be negative log-likelihood or the log score, then this divergence is equivalent to the Kullback-Leibler divergence.
@@
$S_{E}^{(\beta)}(P_{\theta},\mathbf{y})=2\cdot\mathbb{E}||X-\mathbf{y}||^{\beta}-\mathbb{E}||X-X^{\prime}||^{\beta},\quad X,X^{\prime}\sim P_{\theta};\quad\beta\in(0,2)$ (5)
This is a strictly proper scoring rule for the class of probability measures $\mathcal{P}=\{P_{\theta}:\mathbb{E}_{X\sim P_{\theta}}||X||^{\beta}<\infty,\forall\theta\in\Theta\}$, when $\beta\in(0,2)$ (Gneiting and Raftery, 2007). The related divergence is the square of the energy distance, which is a metric between probability distributions. Further, the energy score can be unbiasedly estimated using $\mathbf{x}_{j}\sim P_{\theta},j=1,\ldots,m$ which are independent and identically distributed samples from the model $P_{\theta}$. The unbiased estimate for the energy score in equation (5) can be obtained by Monte Carlo estimates of the expectations in $S_{\mathrm{E}}^{(\beta)}(P,y)$
$\hat{S}_{\mathrm{E}}^{(\beta)}\left(\left\{x_{j}\right\}_{j=1}^{m},\mathbf{y}\right)=\frac{2}{m}\sum_{j=1}^{m}\left\|x_{j}-\mathbf{y}\right\|^{\beta}-\frac{1}{m(m-1)}\sum_{\begin{subarray}{c}j,k=1\\ k\neq j\end{subarray}}^{m}\left\|x_{j}-x_{k}\right\|^{\beta}\quad,\beta\in(0,2).$ (6)
For our implementation, we will consider $\beta=1$ and write $S_{\mathrm{E}}^{(1)}(P,\mathbf{y})$ simply as $S_{\mathrm{E}}(P,\mathbf{y})$.
Theoretical properties of asymptotic normality and generalization bound for scoring rule posteriors have been studied in Giummolè et al. (2019); Pacchiardi (2022); Pacchiardi et al. (2024b). In particular, when the scoring rule is strictly proper, the corresponding posterior contracts around the true model parameter when the model is well-specified. Additionally, the scoring rule posterior for some scoring rules (e.g. energy score or kernel score) exhibits robustness against outliers compared to the standard Bayes posterior using log-score (more details in Chapter 3 of Pacchiardi (2022)).
@@ 3.2 Prequential scoring rule posterior
For complex simulator models–such as deep generative models often used in R; we typically lack closed-form expressions for the conditional distributions, $P_{\theta_{st}}$ and $P_{\theta_{r}}$, which define the log-likelihood of the Markov process (see equation (1)). However, such models can still generate simulations efficiently. Hence combining the idea of scoring rule posterior explained in the previous section with the idea of prequential (predictive sequential) SR proposed in Pacchiardi et al. (2024a), here we propose prequential scoring rule posterior. As before, we use Scoring Rules (SRs) to assess the goodness of fit of the one-step-ahead predictive distribution, conditioned on the previously observed value, to the current observation. By summing a sequence of observations over a period of time, the cumulative SR evaluates the predictive performance of a sequence of conditional models. This cumulative measure is referred to as the prequential score. Thus, on observing the trajectory $H_{T}=h_{T}$, the prequential SR for a parametric MDP ($\mathbb{M}_{\theta}$) based on a scoring rule $S$ can be defined as,
@@
We notice if we have access to the analytical form of $P_{\theta_{st}}$ and $P_{\theta_{r}}$, then taking $S$ to be the log-score the above equation reduces to the log-likelihood defined in equation (1). Moreover, when the scoring rule $S$ is (strictly) proper, the prequential scoring rule $\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T})$ is also (strictly) proper for the class of all Markovian conditional distributions over the next state and reward given the previous state-action pair as shown in Theorem 2 of Pacchiardi et al. (2024a). Also, if we can estimate the SR $S$ as shown in equation (6) unbiasedly, then the prequential SR $\operatorname{\mathcal{PS}}$ can also be estimated unbiasedly by generating $m$ simulations of next states and reward until time $T$, conditioned on previously observed state and action.
Often in RL problems, the reward distribution is chosen deterministically and it is considered as a constant function of state and action. Hence, in those cases, our $\operatorname{\mathcal{PS}}$ defined in equation (7) simplifies to
$\operatorname{\mathcal{PS}}_{T}(\mathbb{M}_{\theta},h_{T})=\sum_{t=1}^{T}S({P}_{\theta}(.|s_{t},a_{t}),s_{t+1}),$ (8)
and $\theta$ only contains parameters $\theta_{st}$. Now suppose $\mathbb{M}_{0}$ is the true data generating process, then we can define the expected prequential score as,
$\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{\theta},\mathbb{M}_{0})=\mathbb{E}_{H_{T}\sim\mathbb{M}_{0}}\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T}).$
From now on, we use the notations $\operatorname{\mathcal{PS}}_{T}(\theta)$ and $\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)$ to denote the empirical prequential score $\operatorname{\mathcal{PS}}_{T}(\mathbb{M}_{\theta},h_{T})$ and the expected prequential score $\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{\theta},\mathbb{M}_{0})$, respectively. Now without loss of generality, we redefine our prequential scoring rule posterior on the parameter $\theta$ of the parametric MDP as,
$\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid h_{T}\right)\propto\pi(\theta)\exp\left\{-w\operatorname{\mathcal{PS}}_{T}(\theta)\right\}.$ (9)
[…]
Asymptotic properties of the generalized prequential posterior:
@@
In this section we derive the asymptotic properties of the generalized prequential posterior. We will prove a Bernstein–von Mises (BvM) theorem under some assumptions. Before stating the assumptions, we first clarify some mathematical notations. We denote the parameter space as $\Theta\subseteq\mathbb{R}^{p}$. We denote the gradient, the matrix of second-order derivatives, and the vector of third-order derivatives of a function $f(\theta)$ with respect to $\theta$ as $f^{\prime}(\theta)=\left(\frac{\partial f(\theta)}{\partial\theta^{i}}(\theta)\right)_{i=1}^{p}\in\mathbb{R}^{p}$, $f^{\prime\prime}(\theta)=\left(\frac{\partial f}{\partial\theta^{i}\partial\theta^{j}}(\theta)\right)_{i,j=1}^{p}\in\mathbb{R}^{p\times p}$ and $f^{\prime\prime\prime}(\theta)=\left(\frac{\partial f}{\partial\theta^{i}\partial\theta^{j}\partial\theta_{k}}(\theta)\right)_{i,j,k=1}^{p}\in\mathbb{R}^{p^{3}}$ respectively. For a given $\theta\in\mathbb{R}^{p}$ and $r>0$, we would denote the open ball of radius $r$ around $\theta$ as $B_{r}(\theta)=\{\theta^{\prime}\in\mathbb{R}^{p}:||\theta^{\prime}-\theta||<r\}$, where |||||$\cdot$||| | ⋅ | | stands for Euclidean norm.
We begin by analyzing the limiting behaviour of the expected prequential score as we observe longer trajectories of interaction with the environment under the following assumption.